OpenStudy (anonymous):

express the perimeter p of a rectangle as a function of length L if the width of the rectangle is twice its length

OpenStudy (anonymous):

p=2(l+b) b=2l p=2(l+2l)=6l:)

OpenStudy (anonymous):

what if the length is twice?

OpenStudy (radar):

Let length = L given, that width W = 2L Perimeter is the sum of the four sides P= L + L + W + W then P= L + L + 2L + 2L P = 6L

OpenStudy (radar):

If it is given that Length is twice the width, then we can say Lengthj equals 2 Width of express it symbolically as L = 2W

OpenStudy (mayankdevnani):

@eli123 w=2l where w=width and l=lenght got it

OpenStudy (radar):

* of to or

OpenStudy (mayankdevnani):

@eli123 satisfy

OpenStudy (mayankdevnani):

ok move towards

OpenStudy (mayankdevnani):

@eli123 hello are you there?